7. Manila, the most populous city, has approximately 4.12x10^7 people


living in it. The population of the entire National Capital Region (NCR) is


approximately 3.1X10^8 people. About how many times greater is the


population of NCR than the population of Manila?


8. A box contains 5.125X10^4 thumb tucks. The mass of each thumb tuck is


8.1X10^-6 kilograms. What is the combined mass of the thumb tucks in


the box?


9. A microscope is set so it makes an object appear 4.3X10^3 times larger


than its usual size. A virus has a diameter of 2.7X10^-8 meter. How large


will the diameter of the virus appear when it is viewed under the


microscope?


10. 1.2 yottameter to yoctometer


1
Expert's answer
2022-04-29T08:59:42-0400

7.

"\dfrac{3.1\times10^8}{4.12\times10^7}\approx7.524\ times"

8.


"5.125\times10^4(8.1\times10^{-6}kg)=0.415125kg"

9.


"4.1\times10^3(2.7\times10^{-8}m)=1.107\times10^{-4}m"

10.

1 yottameter = "10^{48}" yoctometer

1.2 yottameter = "1.2\times10^{48}" yoctometer